# Anomalous Coherence Length in Superconductors with Quantum Metric

### v5

@@ III.1 Derivation on the total coherence length
$\mathcal{M}_{\mathrm{con}}(\bm{q})=-q^{2}v_{F}^{2}f_{4}(\Delta,\mu),$ (S61)
@@
$=\rho_{0}(\epsilon_{F})[\frac{(\mu_{0}-\mu+W/2)(2\Delta^{2}+(\mu_{0}-\mu+W/2)^{2})}{4\Delta^{2}[\Delta^{2}+(\mu_{0}-\mu+W/2)^{2}]^{\frac{3}{2}}}+\frac{(\mu-\mu_{0}+W/2)(2\Delta^{2}+(\mu-\mu_{0}+W/2)^{2})}{4\Delta^{2}[\Delta^{2}+(\mu-\mu_{0}+W/2)^{2}]^{\frac{3}{2}}}].$ (S62)
@@
$\xi=\sqrt{\sqrt{\det\overline{\mathcal{G}}_{ab}}+v_{F}^{2}\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}},$ (S63)
where the factor $\mathcal{M}_{0}(\Delta,\mu)$ and $f_{4}(\Delta,\mu)$ are defined in Eqs. (S54) and (S62).
@@
$\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{1}{4\Delta^{2}}\frac{\frac{(\mu_{0}-\mu+W/2)(2\Delta^{2}+(\mu_{0}-\mu+W/2)^{2})}{[\Delta^{2}+(\mu_{0}-\mu+W/2)^{2}]^{\frac{3}{2}}}+\frac{(\mu-\mu_{0}+W/2)(2\Delta^{2}+(\mu-\mu_{0}+W/2)^{2})}{[\Delta^{2}+(\mu-\mu_{0}+W/2)^{2}]^{\frac{3}{2}}}}{\mathrm{arctanh}\left(\frac{W/2+\mu_{0}-\mu}{\sqrt{\Delta^{2}+(W/2+\mu_{0}-\mu)^{2}}}\right)+\mathrm{arctanh}\left(\frac{W/2-\mu_{0}+\mu}{\sqrt{\Delta^{2}+(W/2-\mu_{0}+\mu)^{2}}}\right)}.$ (S64)
@@
$\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{1}{\Delta^{2}}$ (S65)
@@
$\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{1}{4\Delta^{2}}\frac{1}{\mathrm{arctanh}(1-2\Delta^{2}/\mu_{0}^{2})},$ (S66)
@@
$\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{1}{\Delta^{2}}+\mathcal{O}\left(\frac{W}{\mu_{0}}\right),$ (S67)
@@
$v_{F}^{2}\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\begin{cases}\frac{\text{$v_{F}^{2}$}}{\Delta^{2}}&\frac{W}{\mu_{0}}\rightarrow 1,\frac{\Delta}{\mu_{0}}\gg 1\\ \frac{v_{F}^{2}}{\Delta^{2}}\frac{1}{4\mathrm{arctanh}(1-2\Delta^{2}/\mu_{0}^{2})}&\frac{W}{\mu_{0}}\rightarrow 1,\frac{\Delta}{\mu_{0}}\ll 1\\ \frac{v_{F}^{2}}{\Delta^{2}}&\frac{W}{\mu_{0}}\ll 1\end{cases}$ (S68)
The conventional BCS superconductor corresponds to the case $\frac{W}{\mu_{0}}\ll 1$, where $\sqrt{\det\overline{\mathcal{G}}_{ab}}$ can be suppressed from Eq. (S59).
III.2 Angle dependence of coherence length
@@
$\ell_{qm}(\theta)=\sqrt{\cos^{2}\theta\overline{\mathcal{G}}_{xx}+\sin^{2}\theta\overline{\mathcal{G}}_{yy}+\sin 2\theta\overline{\mathcal{G}}_{xy}}$ (S70)
@@
IV Flat-band lattice model
@@
IV.1 Sawtooth lattice model
In this subsection we study the 1D sawtooth lattice model that host flat band as illustrated in Fig.S1 (a). The unit cell has $A$ and $B$ sites with the hopping terms $V_{1}$ and $V_{2}$. To realize the flat band, $V_{1}=1$ and $V_{2}=\sqrt{2}$. The Hamiltonian in momentum space can be written as
$\mathcal{H}_{\bm{k}}=\left(\begin{matrix}{}-2V_{1}\cos ka&-V_{2}[e^{-ikx}+e^{ik(1-x)}]\\ -V_{2}[e^{ikx}+e^{-ik(1-x)}]&0\end{matrix}\right)~{}.$ (S71)
@@ IV.1 Sawtooth lattice model
$\mathcal{G}_{xx}=\frac{a^{2}+2x(x-1)(1+\cos ka)}{2(2+\cos k)^{2}}~{},$ (S72)
which reaches minimal at $x=0.5a$. The band structure is shown in Fig.S1 (b) with the exact flat band at $E=2$. In the mean-field calculations, we use the attractive interaction $U=0.4$, and we have $\Delta_{A}=0.118$ and $\Delta_{B}=0.088$. The calculated coherence length $\xi$ from Eq. (I.1) is shown in Fig.S1 (c) which is independent of the orbital position. The quantum metric $\sqrt{\overline{\mathcal{G}}_{xx}}$ depends on the orbital position which is parameterized by $x$. $\sqrt{\overline{\mathcal{G}}_{xx}}$ reaches its minimum at $x=0.5$, which is the minimal quantum metric. The discrepancy between the numerical result and the minimal quantum metric can be attributed to the approximation in Eq. (II.2). One can expand the orbital-resolved form factor as $\sum_{\bm{k}\alpha}|u_{\alpha}^{*}(\bm{k}+\bm{q})u_{\alpha}^{*}(\bm{k})|^{2}=\sum_{\bm{k}\alpha}|u_{\alpha}^{*}(\bm{k})|^{4}-\sum_{\bm{k}}\mathcal{G}_{ab}^{o}q_{a}q_{b}$, where $\mathcal{G}_{ab}^{o}$ is a quantity related to quantum metric but not equal to it. The numerical value of the coherence length is consistent with $\ell_{\mathrm{qm}}=\sqrt[4]{\det\overline{\mathcal{G}}^{o}}$, where $\overline{\mathcal{G}}_{ab}^{o}=\sum_{\bm{k}}\mathcal{G}_{ab}^{o}(\bm{k})/\sum_{\bm{k}\alpha}|u_{\alpha}^{*}(\bm{k})|^{4}$. In the following, we will provide the calculations for 2D Lieb lattice, which has similar results.
Figure S1: (a) The sawtooth lattice with $A$ and $B$ atoms forming a unit cell. The $AA$ and $AB$ hoppings are denoted as $V_{1}$ and $V_{2}$, respectively. $x$ is the position of $B$ atoms. (b) The band structure of sawtooth lattice for $V_{1}=1$ and $V_{2}=\sqrt{2}$. (c) The calculated coherence length as a function of $x$ (blue line) and the quantum metric $\sqrt{\overline{\mathcal{G}}_{xx}}$ (orange line). We set the on-site attractive interaction $U=0.4$.
IV.2 Lieb lattice
@@
$\mathcal{H}_{\bm{k}}=\left(\begin{matrix}{}&0&f_{x}&f_{2}\\ &f^{*}_{x}&0&f_{y}\\ &f^{*}_{2}&f^{*}_{y}&0\end{matrix}\right).$ (S73)
@@
To support the analytical results as discussed above, we utilized self-consistent mean-field calculations on the Lieb lattice model, illustrated in Fig.S2 (a), exhibiting a perfectly flat band at $E=0$, as shown in Fig.S2 (b). Each unit cell comprises three atomic orbitals identified as $A$, $B$, and $C$. The staggered hoppings between $A$ and $B$ orbitals are denoted as $(1+\eta)J$ and $(1-\eta)J$, respectively. The quantum metric and band gap can be tuned by varying $\eta$. The next nearest-neighbor hopping is parameterized as $t_{2}$, which induces band dispersion. In Fig.S2 (c), the distribution of $\mathrm{Tr}[\mathcal{G}(\bm{k})]$ at $\eta=0.3$ and $t_{2}=0$ is plotted, reaching a maximum near $M$. We then proceed with fully self-consistent solutions of the on-site order parameter $\hat{\Delta}=\mathrm{diag}(\Delta_{A},\Delta_{B},\Delta_{C})$. The stable solutions indicate $\Delta_{B}\ll\Delta_{A}=\Delta_{C}$ due to the fact that the wave function of the flat band has zero orbital weight on the $B$ site. The intra-unit-cell positions of the orbitals can by varied by $x$, which is the distance between $B$ and $A$/$C$ orbitals.
@@
We numerically evaluate the coherence length $\xi$ by using the multi-orbital correlation function in Eq. (I.1). In Fig. S2(d) we plot $\xi$ as a function of $x$, which reveals that $\xi$ does not depend on the orbital position choice. However, the quantum metric of the flat band is lattice geometry dependent, which reaches the minimum at some specific $x$ corresponding to the minimal quantum metric Huhtinen et al. (2022b). The discrepancy between the numerical result and the minimal quantum metric can be attributed to the approximation in Eq. (II.2). Due to the absence of the dispersion of the bands, the coherence length $\xi$ depends solely on the quantum metric by tuning $\eta$. This is illustrated in Fig. S2(e), where the numerical results of pair correlation functions align with the theoretical results obtained from minimal quantum metric.
@@
To incorporate finite band dispersion, one can introduce an additional nearest-hopping $t_{2}$. This term gives rise to a band dispersion as well as the conventional contribution $\xi_{\mathrm{BCS}}$ to the total coherence length $\xi$. In Fig. S2(f), the total coherence length gradually decreases for $t_{2}=0.01J,0.02J$ when the attractive interaction strength $U$ increases. Remarkably, the total coherence length $\xi$ approaches the flat-band limit due to the suppression of the conventional contribution $\xi_{\mathrm{BCS}}$.
@@
Figure S2: (a) A schematic plot of the Lieb lattice with $A$, $B$ and $C$ orbitals forming a unit cell. The thick and thin lines denote the nearest hoppings with $(1\pm\eta)J$, respectively. A next-nearest hopping $t_{2}$ can make the band dispersive. (b) The calculated band structure of the lattice model at $\eta=0.3$ and $t_{2}=0$ The flat band is separated from the other bands by a gap $\Delta_{\mathrm{gap}}=\sqrt{8}\eta J$. (c) The profiles of quantum metric $\mathrm{Tr}[\mathcal{G}]$ of the flat band in the first Brillouin zone. (d) The calculated coherence length $\xi$ as a function of the orbital position $x$. (e) The quantum metric dependence of $\xi$ as the parameter $\eta$ varies when $U=0.4J$. The red circled line denotes the value from the minimal quantum metric. (f) The calculated $\xi$ at $\eta=0.3$ as a function of the attractive interaction $U/J$ for $t_{2}=0,0.01J,0.02J$. All calculations are conducted at $k_{B}T=0.005J$ and at half-filling $\mu=0$.
IV.3 Flat-band model with tunable quantum metric
@@
In this subsection we would like to study an exact flat-band model with tunable quantum metric and zero Berry curvature. The topologically trivial flat-band model Hamiltonian in Ref. Hofmann et al. (2023b) is given by
$h_{s}(\bm{k})=-t[\sin(\alpha_{\bm{k}})\lambda_{x}+s_{z}\cos(\alpha_{\bm{k}})\lambda_{y}]+[-2t_{2}(\cos k_{x}+\cos k_{y})-\mu]\lambda_{0}~{},$ (S74)
@@
$|u_{+}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ is_{z}e^{is_{z}\alpha_{\bm{k}}}\end{pmatrix},\\ \\ |u_{-}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}-1\\ is_{z}e^{is_{z}\alpha_{\bm{k}}}\end{pmatrix}~{}.$ (S75)
@@
$\mathfrak{G}_{ab}=\frac{1}{4}\partial_{a}\alpha_{\bm{k}}\partial_{b}\alpha_{\bm{k}}.$ (S76)
@@
IV.4 Topological flat-band model with $C=2$
@@
$-\frac{\sqrt{2}}{2}[\cos(k_{x})+\cos(k_{y})]\lambda_{x}-\frac{\sqrt{2}}{2}[-\cos(k_{x})+\cos(k_{y})]\lambda_{y}+\delta_{E}\lambda_{0}.$ (S78)
@@
$\mathcal{G}_{xx}$ $=$ $\frac{2\sin^{2}k_{y}(1+\cos^{2}k_{x}\cos^{2}k_{y})+\cos^{2}k_{y}\sin^{2}k_{x}}{(3+\cos 2k_{x}\cos 2k_{y})^{2}},$ (S79)
$\mathcal{G}_{yy}$ $=$ $\frac{2\sin^{2}k_{x}(1+\cos^{2}k_{x}\cos^{2}k_{y})+\cos^{2}k_{x}\sin^{2}k_{y}}{(3+\cos 2k_{x}\cos 2k_{y})^{2}},$ (S80)
$\mathcal{G}_{xy}$ $=$ $\frac{3\sin 2k_{x}\sin 2k_{y}}{4(3+\cos 2k_{x}\cos 2k_{y})^{2}}.$ (S81)
@@
$\mathcal{H}_{\mathrm{BdG}}=\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}\left[\begin{matrix}{}&h_{\uparrow}(\bm{k})-\mu&\text{diag}(\Delta_{A},\Delta_{B})\\ &\text{diag}(\Delta_{A}^{*},\Delta_{B}^{*})&-h_{\downarrow}(-\bm{k})+\mu\end{matrix}\right]\Psi_{\bm{k}}.$ (S83)
@@
We summarize the flat-band models studied in our work, focusing on the interplay between the quantum metric and band dispersion effects. We provide a general expression for the coherence length as $\xi=\sqrt{\xi_{\mathrm{BCS}}^{2}+\ell_{\mathrm{qm}}^{2}}$ is directly related to the minimal quantum metric. However, we observe a slight discrepancy between the minimal quantum metric and the numerical results from self-consistent mean-field calculations in certain lattice models. This discrepancy is attributed to the uniform pairing assumption in Eq.(II.2). In the main text, we use the band-projection method to study the interplay between the quantum metric and band dispersion effects, and in both cases, the quantum metric is the minimal one. Generally, the minimal quantum metric is a reliable measure for evaluating the coherence length contributed by the quantum metric effect, as seen in twisted moiré materials.
V Continuum model of moiré graphene family
V.1 Continuum model of twisted bilayer graphene
In this section, we provide the calculations of the band structure as well as the quantum metric for twisted bilayer graphene (TBG), twisted trilayer graphene (TTG), and twisted double bilayer graphene (TDBG) in the main text. TBG is formed by two layers with a twist angle $-\theta/2$ and $\theta/2$ relative to $x$ axis respectively (see Fig. S3). In the continuum limit by neglecting the intervalley mixing, the Hamiltonian reads:
$H=H_{1}+H_{2}+H_{int}.$ (S84)
@@ V.1 Continuum model of twisted bilayer graphene
$H_{1}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{1,\xi,\bm{k}}\hbar v_{F}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})a_{1,\xi,\bm{k}},$ (S85)
$H_{2}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{2,\xi,\bm{k}}\hbar v_{F}\hat{R}_{-\theta/2}(\bm{k}-\bm{K_{2}})\cdot(\xi\sigma_{x},\sigma_{y})a_{2,\xi,\bm{k}},$ (S86)
@@
$T(\bm{r})=\left(\begin{array}[]{cc}w_{0}&w_{1}\\ w_{1}&w_{0}\\ \end{array}\right)+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{-i\frac{2\pi}{3}}\\ w_{1}e^{i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)e^{-i\bm{G}_{1}^{M}\cdot\bm{r}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{i\frac{2\pi}{3}}\\ w_{1}e^{-i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)e^{-i(\bm{G}_{1}^{M}+\bm{G}_{2}^{M})\cdot\bm{r}}~{}.$ (S88)

### v6

@@ VIII Competing Interests
$\mathcal{M}_{\mathrm{con}}(\bm{q})=-q^{2}v_{F}^{2}f_{4}(\Delta,\mu),$ (63)
@@
$=\rho_{0}(\epsilon_{F})[\frac{(\mu_{0}-\mu+W/2)(2\Delta^{2}+(\mu_{0}-\mu+W/2)^{2})}{4\Delta^{2}[\Delta^{2}+(\mu_{0}-\mu+W/2)^{2}]^{\frac{3}{2}}}+\frac{(\mu-\mu_{0}+W/2)(2\Delta^{2}+(\mu-\mu_{0}+W/2)^{2})}{4\Delta^{2}[\Delta^{2}+(\mu-\mu_{0}+W/2)^{2}]^{\frac{3}{2}}}].$ (64)
@@
$\xi=\sqrt{\sqrt{\det\overline{\mathcal{G}}_{ab}}+v_{F}^{2}\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}},$ (65)
where the factor $\mathcal{M}_{0}(\Delta,\mu)$ and $f_{4}(\Delta,\mu)$ are defined in Eqs. (56) and (64).
@@
$\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{1}{4\Delta^{2}}\frac{\frac{(\mu_{0}-\mu+W/2)(2\Delta^{2}+(\mu_{0}-\mu+W/2)^{2})}{[\Delta^{2}+(\mu_{0}-\mu+W/2)^{2}]^{\frac{3}{2}}}+\frac{(\mu-\mu_{0}+W/2)(2\Delta^{2}+(\mu-\mu_{0}+W/2)^{2})}{[\Delta^{2}+(\mu-\mu_{0}+W/2)^{2}]^{\frac{3}{2}}}}{\mathrm{arctanh}\left(\frac{W/2+\mu_{0}-\mu}{\sqrt{\Delta^{2}+(W/2+\mu_{0}-\mu)^{2}}}\right)+\mathrm{arctanh}\left(\frac{W/2-\mu_{0}+\mu}{\sqrt{\Delta^{2}+(W/2-\mu_{0}+\mu)^{2}}}\right)}.$ (66)
@@
$\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{1}{\Delta^{2}}$ (67)
@@
$\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{1}{4\Delta^{2}}\frac{1}{\mathrm{arctanh}(1-2\Delta^{2}/\mu_{0}^{2})},$ (68)
@@
$\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\frac{1}{\Delta^{2}}+\mathcal{O}\left(\frac{W}{\mu_{0}}\right),$ (69)
@@
$v_{F}^{2}\frac{f_{4}(\Delta,\mu)}{\mathcal{M}_{0}(\Delta,\mu)}=\begin{cases}\frac{\text{$v_{F}^{2}$}}{\Delta^{2}}&\frac{W}{\mu_{0}}\rightarrow 1,\frac{\Delta}{\mu_{0}}\gg 1\\ \frac{v_{F}^{2}}{\Delta^{2}}\frac{1}{4\mathrm{arctanh}(1-2\Delta^{2}/\mu_{0}^{2})}&\frac{W}{\mu_{0}}\rightarrow 1,\frac{\Delta}{\mu_{0}}\ll 1\\ \frac{v_{F}^{2}}{\Delta^{2}}&\frac{W}{\mu_{0}}\ll 1\end{cases}$ (70)
The conventional BCS superconductor corresponds to the case $\frac{W}{\mu_{0}}\ll 1$, where $\sqrt{\det\overline{\mathcal{G}}_{ab}}$ can be suppressed from Eq. (61).
Angle dependence of coherence length
@@
$\ell_{qm}(\theta)=\sqrt{\cos^{2}\theta\overline{\mathcal{G}}_{xx}+\sin^{2}\theta\overline{\mathcal{G}}_{yy}+\sin 2\theta\overline{\mathcal{G}}_{xy}}$ (72)
@@
Supplementary Note 4: Flat-band lattice model
@@
Sawtooth lattice model
In this subsection we study the 1D sawtooth lattice model that host flat band as illustrated in Supplementary Fig.1 (a). The unit cell has $A$ and $B$ sites with the hopping terms $V_{1}$ and $V_{2}$. To realize the flat band, $V_{1}=1$ and $V_{2}=\sqrt{2}$. The Hamiltonian in momentum space can be written as
$\mathcal{H}_{\bm{k}}=\left(\begin{matrix}{}-2V_{1}\cos ka&-V_{2}[e^{-ikx}+e^{ik(1-x)}]\\ -V_{2}[e^{ikx}+e^{-ik(1-x)}]&0\end{matrix}\right)~{}.$ (73)
@@
$\mathcal{G}_{xx}=\frac{a^{2}+2x(x-1)(1+\cos ka)}{2(2+\cos k)^{2}}~{},$ (74)
which reaches minimal at $x=0.5a$. The band structure is shown in Supplementary Fig.1 (b) with the exact flat band at $E=2$. In the mean-field calculations, we use the attractive interaction $U=0.4$, and we have $\Delta_{A}=0.118$ and $\Delta_{B}=0.088$. The calculated coherence length $\xi$ from Eq. (Mean-field theory for a multi-band superconductor) is shown in Supplementary Fig.1 (c) which is independent of the orbital position. The quantum metric $\sqrt{\overline{\mathcal{G}}_{xx}}$ depends on the orbital position which is parameterized by $x$. $\sqrt{\overline{\mathcal{G}}_{xx}}$ reaches its minimum at $x=0.5$, which is the minimal quantum metric. The discrepancy between the numerical result and the minimal quantum metric can be attributed to the approximation in Eq. (Uniform pairing condition and minimum quantum metric). One can expand the orbital-resolved form factor as $\sum_{\bm{k}\alpha}|u_{\alpha}^{*}(\bm{k}+\bm{q})u_{\alpha}^{*}(\bm{k})|^{2}=\sum_{\bm{k}\alpha}|u_{\alpha}^{*}(\bm{k})|^{4}-\sum_{\bm{k}}\mathcal{G}_{ab}^{o}q_{a}q_{b}$, where $\mathcal{G}_{ab}^{o}$ is a quantity related to quantum metric but not equal to it. The numerical value of the coherence length is consistent with $\ell_{\mathrm{qm}}=\sqrt[4]{\det\overline{\mathcal{G}}^{o}}$, where $\overline{\mathcal{G}}_{ab}^{o}=\sum_{\bm{k}}\mathcal{G}_{ab}^{o}(\bm{k})/\sum_{\bm{k}\alpha}|u_{\alpha}^{*}(\bm{k})|^{4}$. In the following, we will provide the calculations for 2D Lieb lattice, which has similar results.
Supplementary Fig. 1: (a) The sawtooth lattice with $A$ and $B$ atoms forming a unit cell. The $AA$ and $AB$ hoppings are denoted as $V_{1}$ and $V_{2}$, respectively. $x$ is the position of $B$ atoms. (b) The band structure of sawtooth lattice for $V_{1}=1$ and $V_{2}=\sqrt{2}$. (c) The calculated coherence length as a function of $x$ (blue line) and the quantum metric $\sqrt{\overline{\mathcal{G}}_{xx}}$ (orange line). We set the on-site attractive interaction $U=0.4$.
Lieb lattice
@@
$\mathcal{H}_{\bm{k}}=\left(\begin{matrix}{}&0&f_{x}&f_{2}\\ &f^{*}_{x}&0&f_{y}\\ &f^{*}_{2}&f^{*}_{y}&0\end{matrix}\right).$ (75)
@@
To support the analytical results as discussed above, we utilized self-consistent mean-field calculations on the Lieb lattice model, illustrated in Supplementary Fig.2 (a), exhibiting a perfectly flat band at $E=0$, as shown in Supplementary Fig.2 (b). Each unit cell comprises three atomic orbitals identified as $A$, $B$, and $C$. The staggered hoppings between $A$ and $B$ orbitals are denoted as $(1+\eta)J$ and $(1-\eta)J$, respectively. The quantum metric and band gap can be tuned by varying $\eta$. The next nearest-neighbor hopping is parameterized as $t_{2}$, which induces band dispersion. In Supplementary Fig.2 (c), the distribution of $\mathrm{Tr}[\mathcal{G}(\bm{k})]$ at $\eta=0.3$ and $t_{2}=0$ is plotted, reaching a maximum near $M$. We then proceed with fully self-consistent solutions of the on-site order parameter $\hat{\Delta}=\mathrm{diag}(\Delta_{A},\Delta_{B},\Delta_{C})$. The stable solutions indicate $\Delta_{B}\ll\Delta_{A}=\Delta_{C}$ due to the fact that the wave function of the flat band has zero orbital weight on the $B$ site. The intra-unit-cell positions of the orbitals can by varied by $x$, which is the distance between $B$ and $A$/$C$ orbitals.
@@
We numerically evaluate the coherence length $\xi$ by using the multi-orbital correlation function in Eq. (Mean-field theory for a multi-band superconductor). In Supplementary Fig. 2(d) we plot $\xi$ as a function of $x$, which reveals that $\xi$ does not depend on the orbital position choice. However, the quantum metric of the flat band is lattice geometry dependent, which reaches the minimum at some specific $x$ corresponding to the minimal quantum metric Huhtinen et al. (2022). The discrepancy between the numerical result and the minimal quantum metric can be attributed to the approximation in Eq. (Uniform pairing condition and minimum quantum metric). Due to the absence of the dispersion of the bands, the coherence length $\xi$ depends solely on the quantum metric by tuning $\eta$. This is illustrated in Supplementary Fig. 2(e), where the numerical results of pair correlation functions align with the theoretical results obtained from minimal quantum metric.
@@
To incorporate finite band dispersion, one can introduce an additional nearest-hopping $t_{2}$. This term gives rise to a band dispersion as well as the conventional contribution $\xi_{\mathrm{BCS}}$ to the total coherence length $\xi$. In Supplementary Supplementary Fig. 2(f), the total coherence length gradually decreases for $t_{2}=0.01J,0.02J$ when the attractive interaction strength $U$ increases. Remarkably, the total coherence length $\xi$ approaches the flat-band limit due to the suppression of the conventional contribution $\xi_{\mathrm{BCS}}$.
@@
Supplementary Fig. 2: (a) A schematic plot of the Lieb lattice with $A$, $B$ and $C$ orbitals forming a unit cell. The thick and thin lines denote the nearest hoppings with $(1\pm\eta)J$, respectively. A next-nearest hopping $t_{2}$ can make the band dispersive. (b) The calculated band structure of the lattice model at $\eta=0.3$ and $t_{2}=0$ The flat band is separated from the other bands by a gap $\Delta_{\mathrm{gap}}=\sqrt{8}\eta J$. (c) The profiles of quantum metric $\mathrm{Tr}[\mathcal{G}]$ of the flat band in the first Brillouin zone. (d) The calculated coherence length $\xi$ as a function of the orbital position $x$. (e) The quantum metric dependence of $\xi$ as the parameter $\eta$ varies when $U=0.4J$. The red circled line denotes the value from the minimal quantum metric. (f) The calculated $\xi$ at $\eta=0.3$ as a function of the attractive interaction $U/J$ for $t_{2}=0,0.01J,0.02J$. All calculations are conducted at $k_{B}T=0.005J$ and at half-filling $\mu=0$.
Flat-band model with tunable quantum metric
@@
In this subsection we would like to study an exact flat-band model with tunable quantum metric and zero Berry curvature. The topologically trivial flat-band model Hamiltonian in Ref. Hofmann et al. (2023) is given by
$h_{s}(\bm{k})=-t[\sin(\alpha_{\bm{k}})\lambda_{x}+s_{z}\cos(\alpha_{\bm{k}})\lambda_{y}]+[-2t_{2}(\cos k_{x}+\cos k_{y})-\mu]\lambda_{0}~{},$ (76)
@@
$|u_{+}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ is_{z}e^{is_{z}\alpha_{\bm{k}}}\end{pmatrix},\\ \\ |u_{-}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}-1\\ is_{z}e^{is_{z}\alpha_{\bm{k}}}\end{pmatrix}~{}.$ (77)
@@
$\mathfrak{G}_{ab}=\frac{1}{4}\partial_{a}\alpha_{\bm{k}}\partial_{b}\alpha_{\bm{k}}.$ (78)
@@
Topological flat-band model with $C=2$
@@
$-\frac{\sqrt{2}}{2}[\cos(k_{x})+\cos(k_{y})]\lambda_{x}-\frac{\sqrt{2}}{2}[-\cos(k_{x})+\cos(k_{y})]\lambda_{y}+\delta_{E}\lambda_{0}.$ (80)
@@
$\mathcal{G}_{xx}$ $=$ $\frac{2\sin^{2}k_{y}(1+\cos^{2}k_{x}\cos^{2}k_{y})+\cos^{2}k_{y}\sin^{2}k_{x}}{(3+\cos 2k_{x}\cos 2k_{y})^{2}},$ (81)
$\mathcal{G}_{yy}$ $=$ $\frac{2\sin^{2}k_{x}(1+\cos^{2}k_{x}\cos^{2}k_{y})+\cos^{2}k_{x}\sin^{2}k_{y}}{(3+\cos 2k_{x}\cos 2k_{y})^{2}},$ (82)
$\mathcal{G}_{xy}$ $=$ $\frac{3\sin 2k_{x}\sin 2k_{y}}{4(3+\cos 2k_{x}\cos 2k_{y})^{2}}.$ (83)
@@
$\mathcal{H}_{\mathrm{BdG}}=\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}\left[\begin{matrix}{}&h_{\uparrow}(\bm{k})-\mu&\text{diag}(\Delta_{A},\Delta_{B})\\ &\text{diag}(\Delta_{A}^{*},\Delta_{B}^{*})&-h_{\downarrow}(-\bm{k})+\mu\end{matrix}\right]\Psi_{\bm{k}}.$ (85)
@@
We summarize the flat-band models studied in our work, focusing on the interplay between the quantum metric and band dispersion effects. We provide a general expression for the coherence length as $\xi=\sqrt{\xi_{\mathrm{BCS}}^{2}+\ell_{\mathrm{qm}}^{2}}$ is directly related to the minimal quantum metric. However, we observe a slight discrepancy between the minimal quantum metric and the numerical results from self-consistent mean-field calculations in certain lattice models. This discrepancy is attributed to the uniform pairing assumption in Eq.(Uniform pairing condition and minimum quantum metric). In the main text, we use the band-projection method to study the interplay between the quantum metric and band dispersion effects, and in both cases, the quantum metric is the minimal one. Generally, the minimal quantum metric is a reliable measure for evaluating the coherence length contributed by the quantum metric effect, as seen in twisted moiré materials.
Supplementary Note 5: Continuum model of moiré graphene family
Continuum model of twisted bilayer graphene
In this section, we provide the calculations of the band structure as well as the quantum metric for twisted bilayer graphene (TBG), twisted trilayer graphene (TTG), and twisted double bilayer graphene (TDBG) in the main text. TBG is formed by two layers with a twist angle $-\theta/2$ and $\theta/2$ relative to $x$ axis respectively (see Supplementary Fig. 3). In the continuum limit by neglecting the intervalley mixing, the Hamiltonian reads:
$H=H_{1}+H_{2}+H_{int}.$ (86)
@@
$H_{1}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{1,\xi,\bm{k}}\hbar v_{F}\hat{R}_{\theta/2}(\bm{k}-\bm{K_{1}})\cdot(\xi\sigma_{x},\sigma_{y})a_{1,\xi,\bm{k}},$ (87)
$H_{2}$ $=\sum_{\xi,\bm{k}}a^{\dagger}_{2,\xi,\bm{k}}\hbar v_{F}\hat{R}_{-\theta/2}(\bm{k}-\bm{K_{2}})\cdot(\xi\sigma_{x},\sigma_{y})a_{2,\xi,\bm{k}},$ (88)
@@
$T(\bm{r})=\left(\begin{array}[]{cc}w_{0}&w_{1}\\ w_{1}&w_{0}\\ \end{array}\right)+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{-i\frac{2\pi}{3}}\\ w_{1}e^{i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)e^{-i\bm{G}_{1}^{M}\cdot\bm{r}}+\left(\begin{array}[]{cc}w_{0}&w_{1}e^{i\frac{2\pi}{3}}\\ w_{1}e^{-i\frac{2\pi}{3}}&w_{0}\\ \end{array}\right)e^{-i(\bm{G}_{1}^{M}+\bm{G}_{2}^{M})\cdot\bm{r}}~{}.$ (90)